OpenStudy (anonymous):

Tony Saulino recently refinanced his $265,000 home from a 7.5%, 30-year mortgage to a 5%,15-year mortgage. alculate the monthly payment for both orgages. Then calculate an estimate of how much interest he is savingwith the 15-year mortgage.

OpenStudy (tkhunny):

1) At the beginning of the loan? That's an odd time to refinance. 2) A question about interest, comparing two different periods, is a very odd question. One must somehow make sense of the 15 years after the 15-year loan is taken care of. v30 = 1/(1 + 0.075/12) v15 = 1/(1 + 0.050/12) \(265000 = P30\cdot v30\cdot\dfrac{1-v30^{360}}{1-v30}\) \(265000 = P15\cdot v15\cdot\dfrac{1-v15^{180}}{1-v15}\) \(360*P30 - 180*P15\) Solves the problem. Understanding the basic principles solves all such problems.

OpenStudy (anonymous):

Thank You that helped me alot